anyone help me those questions? thx!!Question 1 dz , + )2, where a1, a2, and a3 are constants. Find dq. Not yet answered Let-(o)-1(a Points out of 1 Select one Flag question a. 3q Question Not yet answered Points out of 1 Let w(x,y) b1 + b2x + bay + b,xy, where x and y are variables and b, are all constants. Find ay. Select one Flag question a. ay C. ay d. ayQuestion 3 Suppose your grade on a test (g) is a function of hours spent studying (s) studying an extra hour on your test grade? Select one: Not yet answered g(S) 10 log (s). Which of the following represents the marginal effect of Points out of 1 Flag question dg 10 a. ds dg-10 ds 10 C. dg Question 4 Not yet answered Points out of 1 Suppose that a workers log-wage is determined solely by a variable representing years of education (Edu), its square, and constants a, b, and c such that logw-a+ b Edu + c Edu2. What is the correct interpretation of the marginal effect of education on log-wage? Hint: this is asking about dElogw) Flag question dw not dEdu Select one: a. A one year increase in education increases wage by (b-Edu >< 100)% b. A one year increase in education increases wage by (b 100)96 c. A one year increase in education increases wage by ((b + 2c . Edu) 100196 d. A one year increase in education increases wage by (2c . Edu C 100)%The expression (Σ, λΜ), can be expanded to equal which of the following? Question 5 Not yet answered Points out of 1 Select one Flag question 22,2 Question 6 Suppose that y is a random variable with a normal distribution, which can be Not yet answered represented by y N(H,2). Which of the following represents the transformation to a standard normal distribution? Points out of 1 Flag question Select one a N(0.1) b. ~ N(0,1) C.σ2
Answer #1

1. z(q) = (a1 + q)2 + (a2 + q)2 + (a3 + q)2

dz/ dq = 2q + 2 a1 + 2q+ 2 a2 + 2q + 2 a3 = 6q + 2(a1 + a2 + a3)= 2 sum(ai) + 6q .The correct option is b.

2. dw/ dy = b3 + b4 x. The correct option is b.

3. g(s) = 10 log(s)

dg/ ds = 10 /s .The correct option is a.

4. d(logw) / d Edu = b + 2c Edu

The correct option is c.

5. The correct option is b.

6. The correct option is d.